OpenStudy (anonymous):

7) Which equation below is an identity? a) 5f + 6 = 5f – 6 b) 4q + 3 = 3q – 3 + q c) 6t + 1 – 3t = –2t – 1 d) 8a – (a + 4) = 9a – 2(a + 2) 8) Which equation below has no solution? a) 4z – 1 = –4z + 10 b) 7n + 9 = 7n – 2 c) 6s + 5 = 5s – 6 d) 8c – 3 = 8c + 7 – c 9) Chris and Lucy went to a baseball game. They each bought a bag of peanuts and Lucy purchased a bottled water. Chris had a $10 gift card toward the cost and paid the rest, which came to $32.00. If a baseball ticket costs $14.75 and a bag of peanuts costs $4.50, what is the cost of a bottled water?

OpenStudy (anonymous):

solve the equation. 10) –11 = q + 8 11) d – 6 = –13 12) 12.9 = 3m 13) −4= y/7 14) −5=13/42x 15) 3(b + 2) – b = 2b + 6 16) 11 – 4r = 5r – 11 – 9r 17) 4(h – 3) – 9 = 3 18) 72 = –2(–6 + 5y) 19) –4x + 7 = 5(2x – 1) – 3(–12 + 2x) 20) 8= j−4/5 21) 17+z/3=43

OpenStudy (anonymous):

7) you put variable in one side and numbers in another side. After calculating if you get the form 0=0 then it's identity 8) Same as (7) However after calculating the variable is still there, and equal to a constant

OpenStudy (anonymous):

hope it help